# The systemic recoil velocity distribution and the scale height of isolated millisecond pulsar systems: Implications on neutron star retention fractions in star clusters

### v1

@@ Appendix A Selection of the millisecond pulsar sample
PSR $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. $a$$a$$a$References for proper motions and parallaxes: (1) Bangale et al. (2024), (2) Alam et al. (2020), (3) Ding et al. (2023), (4) Shamohammadi et al. (2024), (5) Curyło et al. (2023), (6) Reardon et al. (2024), (7) van der Wateren et al. (2022), (8) Antoniadis et al. (2023b), (9) Guillemot et al. (2016), (10) Fiore et al. (2023), (11) Kramer et al. (2021), (12) Deller et al. (2019), (13) Deller et al. (2012), (14) Jennings et al. (2018), (15) Konacki & Wolszczan (2003), (16) Yan et al. (2013), (17) Swiggum et al. (2017), (18) Reardon et al. (2021), (19) Chatterjee et al. (2009), (20) Arzoumanian et al. (2018), (21) Deller et al. (2018), (22) Geyer et al. (2023), (23) Ding et al. (2024a), (24) Stovall et al. (2019), (25) Alam et al. (2021), and (26) Crowter et al. (2020).
$d$ $z$ $v_{l}$ $v_{b}$ $v_{\perp}$
($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) ($\mathrm{mas\leavevmode\nobreak\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
J1939$+$2134 $0.074(2)$ $-0.410(3)$ $0.35(3)$ (3) $3.0^{+0.3}_{-0.2}$ $-0.015(1)$ $80^{+9}_{-7}$ $-3.8^{+0.3}_{-0.4}$ $80^{+9}_{-8}$
J1946--5403 $-1.08(2)$ $-4.75(5)$ $0.8(1)$ (4) $1.3(2)$ $-0.65^{+0.08}_{-0.11}$ $-18^{+4}_{-5}$ $7^{+2}_{-1}$ $19^{+5}_{-4}$
J2043$+$1711 $-5.70(1)$ $-10.84(2)$ $0.72(7)$ (2) $1.4^{+0.2}_{-0.1}$ $-0.38(4)$ $-43^{+4}_{-5}$ $-13^{+1}_{-2}$ $45^{+5}_{-4}$
J2124--3358 $-14.09(1)$ $-50.32(3)$ $2.1(1)$ (8) $0.48(2)$ $-0.34(2)$ $-106^{+5}_{-6}$ $20.6^{+1.0}_{-0.9}$ $108^{+6}_{-5}$
J2129--0429 $12.10(7)$ $10.19(6)$ $0.48(7)$ (1), (14) $2.1^{+0.4}_{-0.3}$ $-1.3(2)$ $194^{+32}_{-25}$ $-63^{+9}_{-11}$ $204^{+37}_{-27}$
J2145--0750 $-9.7(3)$ $-8.2(9)$ $1.60(6)$ (12) $0.63(2)$ $-0.42(2)$ $-21(3)$ $14(2)$ $25(2)$
J2222--0137 $44.707(5)$ $-5.40(1)$ $3.72(1)$ (23) $0.269(1)$ $-0.1935(7)$ $39.8(1)$ $-43.3(2)$ $58.8(2)$
J2234$+$0611 $25.30(2)$ $9.71(5)$ $1.03(4)$ (24) $0.97(4)$ $-0.66(3)$ $139^{+6}_{-5}$ $-44(2)$ $145^{+6}_{-5}$
J2234$+$0944 $7.0(2)$ $-33.2(5)$ $1.4(3)$ (25) $0.9^{+0.3}_{-0.2}$ $-0.6^{+0.1}_{-0.2}$ $-46^{+11}_{-18}$ $-116^{+26}_{-45}$ $124^{+49}_{-28}$
J2241--5236 $18.87(1)$ $-5.28(1)$ $0.96(4)$ (18) $1.05^{+0.05}_{-0.04}$ $-0.86^{+0.03}_{-0.04}$ $-62(3)$ $-61^{+2}_{-3}$ $87(4)$
J2256--1024 $3(1)$ $-8(2)$ $0.5(2)$ (26) $2.1^{+0.7}_{-0.5}$ $-1.8^{+0.4}_{-0.6}$ $-18^{+17}_{-21}$ $-79^{+22}_{-35}$ $83^{+35}_{-25}$
J2317$+$1439 $-1.46(10)$ $3.7(2)$ $0.60(8)$ (2) $1.7^{+0.3}_{-0.2}$ $-1.2^{+0.1}_{-0.2}$ $40^{+6}_{-5}$ $42^{+6}_{-5}$ $58^{+9}_{-7}$
J2322$+$2057 $-18.30(5)$ $-14.9(1)$ $1.2(3)$ (8), (4) $1.0^{+0.5}_{-0.2}$ $-0.6^{+0.2}_{-0.3}$ $-91^{+22}_{-43}$ $-19^{+6}_{-12}$ $93^{+43}_{-23}$
J2322--2650 $-2.37(9)$ $-8.2(2)$ $1.3(2)$ (4) $0.8^{+0.2}_{-0.1}$ $-0.8^{+0.1}_{-0.2}$ $-21^{+4}_{-6}$ $7.4^{+0.6}_{-0.5}$ $23^{+6}_{-4}$
@@
333 When an MSP has multiple proper motions and parallaxes reported by different studies, the values with smaller uncertainties are adopted. Despite our meticulous efforts, some of the compiled proper motions and parallaxes may not represent the most precise measurements available.
@@ B.1 Candidate probability density functions of $v_{\mathrm{l}}$
[…]
@@
Table 4: Best-fit parameters of candidate $v_{\mathrm{l}}$ probability density functions defined by Equations 2 and 3, along with the Bayes factors $K$ and the goodness differences $\Delta\Lambda$ calculated with the best-fit Cauchy distribution ($N=1$ and $\gamma_{1}=51.8$$\mathrm{km\leavevmode\nobreak\ s^{-1}}$) as the reference of comparison.
$N$ $\gamma_{1}$ $\alpha_{1}$ $\gamma_{2}$ $\alpha_{2}$ $\gamma_{3}$ $\alpha_{3}$ $\gamma_{4}$ $\alpha_{4}$ $\ln{K}$ $\Delta\Lambda$
($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$) ($\mathrm{km\leavevmode\nobreak\ s^{-1}}$)
@@ B.2 Selecting from the candidate probability density distributions
From each of the two prime PDF groups introduced in Appendix B.1, we selected one $v_{\mathrm{l}}$ PDF based on a criterion that is equivalent to the Akaike information criterion (Akaike 2011). Specifically, the goodness $\Lambda$ of a PDF is quantified by
@@
$L=\displaystyle\prod_{j=1}^{65}p\left(v^{j}_{\mathrm{l}}\right)\,\,,$ (6)
where $p\left(v^{j}_{\mathrm{l}}\right)$ is the value of Equation 2 or 3 at the observed $v^{j}_{\mathrm{l}}\,\,(j=1,2,...,65)$.
@@
where $K\equiv L_{1}/L_{2}$ is the Bayes factor between the two PDFs. To evaluate the uncertainty in $K$ due to the uncertainties in the 65 $v_{\mathrm{l}}$ measurements, we assumed the observed $v_{\mathrm{l}}$ follow split normal distributions, and simulated 10000 sets of 65 $v_{\mathrm{l}}$ values. From each set of 65 $v_{\mathrm{l}}$ values we derived a $K$ based on Equation 6. With the altogether 10000 derived $K$ values, we estimated $K$ and its uncertainty for two PDFs of comparison. The calculated $\ln{K}$ (as well as their uncertainties) of the candidate PDFs with respect to the best-fit Cauchy distribution are provided in Table 4.
@@
Between the two finalist PDFs, the Cauchy distribution is preferred over the three-component normal distribution according to the $\Lambda$ standard (see Table 4). Therefore, we conclude that the sample of 65 observed $v_{\mathrm{l}}$ is well described by the simple PDF
$\frac{p}{1\,\mathrm{s\leavevmode\nobreak\ {km}^{-1}}}=\frac{1}{51.8\pi\left[1+\left(\frac{{v_{\mathrm{l}}}}{51.8\,\mathrm{km\leavevmode\nobreak\ s^{-1}}}\right)^{2}\right]}\,\,.$ (8)
@@ C.1 Millisecond pulsar age distribution
The age distribution of MSPs is poorly constrained. We approached the rough MSP age distribution by the following procedure. Firstly, we identified MSPs from the PSRCAT1 catalogue with the MSP criteria (i) and (ii) stated in Appendix A. After removing sources without constraint on characteristic age $\tau_{\mathrm{c}}$, we obtained 399 MSP $\tau_{\mathrm{c}}$ values. To derive the age distribution, we corrected the $\tau_{\mathrm{c}}$ sample by a sample of age correction factors (i.e., ratio between age and $\tau_{\mathrm{c}}$) estimated by Kiziltan & Thorsett (2010) for 42 MSPs having proper motion measurements (i.e., the reciprocals of “$\tau_{c}/\tilde{\tau}^{\prime}$” reported in Table 1 of Kiziltan & Thorsett 2010). Specifically, we multiplied each of the $\tau_{\mathrm{c}}$ sample by each of the correction factor sample, which led to a sample of $399\times 42=16,758$ MSP ages. After discarding ages larger than the age of the Galaxy (i.e., 13.6 Gyr), we obtained a sample of 13,005 MSP ages (see Figure 6 for the histogram of this sample). We considered this age sample as an indicative approximation to the MSP age distribution, and adopted the age sample in the DPS analysis (see Section 4.2).
@@ C.2.1 The underlying probability density function of the observed Galactic heights
When searching for the candidates for the underlying PDF of the observed $z$ sample, we assumed that the PDF is both uni-modal and symmetric about zero. The same assumptions had been adopted in identifying the underlying PDF for the observed $v_{\mathrm{l}}$ (see Section 3). We found that the one-parameter Cauchy distribution and the one-parameter Laplace distribution are both plausible candidates for the underlying PDF of the observed $z$ (see Figure 2). However, further analysis following the method detailed in Appendix B.2 prefers the best-fit Laplace distribution over the Cauchy counterpart with $\ln{K}=8.0^{+0.7}_{-0.9}$, where $K$ is the Bayes factor. The best-fit Laplace distribution for the observed $z$ sample has the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ of 0.52 kpc. We note that $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ is derived from nearby isolated MSP system, which should be distinguished from the scale height $\zeta_{1}$ of Galaxy-wide isolated MSP systems.
@@ C.2.2 The determination of the Galaxy-wide birth scale height $\zeta_{0}$
As mentioned in Section 4.1, various $\zeta_{0}$ would lead to different levels of the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$. Due to the complex distributions of the observed $z$ and $v_{\mathrm{b}}$ (see Figure 2 or Figure 10), we find it difficult to determine the $\zeta_{0}$ using statistical methods (e.g., Kolmogorov-Smirnov tests or the Wasserstein metric) that quantify the similarity between the observed samples and the DPS counterparts. On the other hand, the evolved scale height $\zeta_{\mathrm{e}}^{\mathrm{local}}$ is expected to increase with $\zeta_{0}$ (see Figure 10), and match the observed counterpart $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Appendix C.2.1). Therefore, we reran the DPS with incrementally larger $\zeta_{0}$ until the best-fit $\zeta_{\mathrm{e}}^{\mathrm{local}}$ reaches 0.52 kpc. In this way, we found $\zeta_{0}\approx 0.32$ kpc. The dependence of the evolved $v_{\mathrm{b}}$, $z_{\mathrm{e}}$ and $v_{\mathrm{l}}$ on $\zeta_{0}$ is illustrated in Figure 10. As is shown in Figure 10, both the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$ match the observed counterparts at $\zeta_{0}=0.32$ kpc. Additionally, the average levels of both evolved $|\mathrm{b}|$ and $|z_{\mathrm{e}}|$ rise with $\zeta_{0}$ as expected. In comparison, the evolved $v_{\mathrm{l}}$ stays robust with respect to changes in the birth vertical distribution of isolated MSP systems.
@@
Figure 10: The smoothed histogram profiles of $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ derived from dynamical population synthesis based on different Galaxy-wide birth scale height $\zeta_{0}$ of isolated MSP systems. Overlaid are the visualized histograms (and their smoothed profiles) of the observed $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ (see Section 2).
@@
Finally, we note that the observed $z$ sample might be biased by Galactic-latitude-related selection effects. As a major example, pulsars located at low Galactic latitudes ($b$) are more susceptible to propagation effects caused by the ionized interstellar medium (IISM), which might reduce the detection rate of these low-$b$ pulsars. This relative drop in the number of low-$b$ pulsars would become increasingly significant as further-away pulsars are searched and observed with large telescopes, but may still have small impact on this work where the vast majority of MSPs are in the vicinity of the Solar system (see Table 2). On the other hand, most pulsar search programs have been focusing on low-$b$ sky regions, which serves as another selection effect that would increase the relative number of low-$b$ pulsars. For this work, we see no need to take into account the impact of complex $b$-related selection effects on the observed $z$, especially as the DPS results based on $\zeta_{0}=0.32$ kpc agree with both the observed $z$ and the observed $v_{\mathrm{b}}$ samples. However, if the observed $z$ and the observed $v_{\mathrm{b}}$ suggest different $\zeta_{0}$, this may indicate that $b$-related selection effects have to be accounted for.
@@
Figure 12: The histogram of Galaxy-wide evolved Galactic heights $z_{\mathrm{e}}$ derived from dynamical population synthesis. Overlaid are the best-fit Laplace and Cauchy distributions of the $z_{\mathrm{e}}$ sample, with the scale parameters of 0.66 kpc and 0.47 kpc, respectively.

### v3

@@ A.1 High-precision samples
PSR solitary? $\mu_{\alpha}$ $\mu_{\delta}$ $\varpi$ Ref. aaaaaaReferences for proper motions and parallaxes: (1) Bangale24, (2) alam20, (3) Ding23, (4) Shamohammadi24, (5) curyo23, (6) reardon24, (7) van-der-Wateren22, (8) Antoniadis23, (9) Guillemot16, (10) fiore23, (11) Kramer21a, (12) Deller19, (13) deller12, (14) Jennings18, (15) Konacki03, (16) Yan13, (17) Swiggum17, (18) reardon21, (19) Chatterjee09, (20) Arzoumanian18a, (21) Deller18, (22) geyer23, (23) ding24c, (24) Stovall19, (25) alam21, and (26) Crowter20.
$d$ $z$ $v_{l}$ $_{b} vv_{\perp}$
(Y/N) ($\mathrm{mas\penalty 10000\ yr^{-1}}$) ($\mathrm{mas\penalty 10000\ yr^{-1}}$) (mas) (kpc) (kpc) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$)
J1946-5403 N 1.08(3)-1.08(3) 4.75(5)-4.75(5) $0.8(1)$ (4) $1.3(2)$ $0.65^{+0.08}_{-0.11}$ $17^{+3}_{-5}$ $7^{+2}_{-1}$ $19^{+5}_{-4}$
J2043++1711 N 5.70(1)-5.70(1) 10.84(2)-10.84(2) $0.72(6)$ (2) $1.4(1)$ $0.38^{+0.03}_{-0.04}$ 42(4)-42(4) 13(1)-13(1) $44^{+5}_{-4}$
J2124-3358 Y 14.09(1)-14.09(1) 50.32(3)-50.32(3) $2.1(1)$ (8) $0.48(2)$ 0.34(2)-0.34(2) 107(6)-107(6) $20.6^{+1.0}_{-0.9}$ $109(6)$
J2129-0429 N $12.10(7)$ $10.19(6)$ $0.48(7)$ (1), (14) $2.1^{+0.4}_{-0.3}$ 1.3(2)-1.3(2) $194^{+33}_{-25}$ $63^{+9}_{-12}$ $204^{+35}_{-27}$
J2145-0750 N 9.7(3)-9.7(3) 8.2(9)-8.2(9) $1.60(6)$ (12) $0.63(2)$ 0.42(2)-0.42(2) 21(3)-21(3) $14(2)$ $25(2)$
J2222-0137 N $44.707(5)$ 5.40(1)-5.40(1) $3.72(1)$ (23) $0.269(1)$ 0.1935(7)-0.1935(7) $39.8(1)$ 43.3(2)-43.3(2) $58.8(2)$
J2234++0611 N $25.30(2)$ $9.71(5)$ $1.03(4)$ (24) $0.97(4)$ 0.66(3)-0.66(3) $139(5)$ 44(2)-44(2) $145^{+6}_{-5}$
J2234++0944 N $6.9(2)$ 33.2(5)-33.2(5) $1.4(3)$ (4), (25) $0.8^{+0.2}_{-0.1}$ $0.52^{+0.10}_{-0.15}$ $43^{+9}_{-13}$ $08^{+21}_{-33} 11$$17^{+36}_{-23}$
J2241-5236 N $18.881(4)$ 5.294(5)-5.294(5) $0.96(4)$ (18) $1.05^{+0.05}_{-0.04}$ $0.86^{+0.03}_{-0.04}$ 62(3)-62(3) 61(2)-61(2) $87(4)$
J2256-1024 N $3(1)$ 8(2)-8(2) $0.5(2)$ (26) $2.1^{+0.7}_{-0.5}$ $1.8^{+0.4}_{-0.6}$ $18^{+18}_{-20}$ $79^{+23}_{-33}$ $83^{+36}_{-25}$
J2317++1439 N 1.46(10)-1.46(10) $3.7(2)$ $0.60(8)$ (2) $1.7^{+0.3}_{-0.2}$ $1.2^{+0.1}_{-0.2}$ $40^{+6}_{-5}$ $42^{+6}_{-5}$ $58^{+9}_{-7}$
J2322++2057 Y 18.30(5)-18.30(5) 14.9(1)-14.9(1) $1.2(3)$ (4), (8) $1.0^{+0.5}_{-0.2}$ $0.6^{+0.2}_{-0.3}$ $91^{+22}_{-42}$ $19^{+6}_{-11}$ $93^{+43}_{-23}$
J2322-2650 N 2.37(9)-2.37(9) 8.2(2)-8.2(2) $1.3(2)$ (4) $0.8^{+0.2}_{-0.1}$ $0.8^{+0.1}_{-0.2}$ $22^{+4}_{-6}$ $7.4^{+0.6}_{-0.5}$ $23^{+6}_{-4}$
@@
 [^5] When an MSP has multiple proper motions and parallaxes reported by different studies, the values with smaller uncertainties (instead of higher significances) are adopted. Despite our meticulous efforts, some of the compiled proper motions and parallaxes may not represent the most precise measurements available.
Figure 8: The upper-left, lower-left and lower-right panels display the normalized histograms of Galactic heights $z$, the Galactic-longitude component $v_{\mathrm{l}}$ and the Galactic-latitude component $v_{\mathrm{b}}$ of transverse peculiar velocities, respectively. The normalized histograms have been smoothed with kernel density estimation (using scipy.stats.Gaussian_kde). Each histogram is concatenated from 10000 simulations drawn from the assumed split normal distributions for the measurements (of $z$, $v_{\mathrm{l}}$ or $v_{\mathrm{b}}$; see Table 6). The black, magenta and yellow curves correspond to the full sample (of 64 field MSPs) and two high-precision samples (see Appendix A.1), respectively.
@@ B.1 Candidate probability density functions of $v_{\mathrm{l}}$
$\cases{p}=\displaystyle\sum_{i}^{N}\frac{\alpha_{i}}{\sqrt{2\pi{\gamma_{i}}^{2}}}\exp{\left(-\frac{{v_{\mathrm{l}}}^{2}}{2{\gamma_{i}}^{2}}\right)}\,\,\,\,\,\,\,\,\,\,(\alpha_{i}>0)\\ \displaystyle\sum_{i}^{N}\alpha_{i}=1\,\,\,.{}$ (3)
@@
Table 8: Best-fit parameters of candidate $v_{\mathrm{l}}$ probability density functions defined by Equations 2 and 3, along with the Bayes factors $K$ and the goodness differences $\Delta\Lambda$ calculated with the best-fit Cauchy distribution ($N=1$ and $\gamma_{1}=51.2$$\mathrm{km\penalty 10000\ s^{-1}}$) as the reference of comparison. [^6] Despite our exhaustive efforts, the best-fit two-component Cauchy distribution closely approaches the one-component Cauchy distribution, as the scales of the two components are nearly identical.
NN $\gamma_{1}$ $\alpha_{1}$ $\gamma_{2}$ $\alpha_{2}$ $\gamma_{3}$ $\alpha_{3}$ $\gamma_{4}$ $\alpha_{4}$ $\ln{K}$ $\Delta\Lambda$
($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$) ($\mathrm{km\penalty 10000\ s^{-1}}$)
@@ B.2 Selecting from the candidate probability density distributions
From each of the two prime PDF groups introduced in Appendix B.1, we selected one $v_{\mathrm{l}}$ PDF based on a criterion that is equivalent to the Akaike information criterion (Akaike11). Specifically, the goodness $\Lambda$ of a PDF is quantified by
@@
$L=\displaystyle\prod_{j=1}^{64}p\left(v^{j}_{\mathrm{l}}\right)\,\,$, (6)
where $p\left(v^{j}_{\mathrm{l}}\right)$ is the value of Equation 2 or 3 at the observed $v^{j}_{\mathrm{l}}\,\,(j=1,2,...,64)$.
@@
where $K\equiv L_{1}/L_{2}$ is the Bayes factor between the two PDFs. To evaluate the uncertainty in $K$ due to the uncertainties in the 64 $v_{\mathrm{l}}$ measurements, we assumed the observed $v_{\mathrm{l}}$ follow split normal distributions, and simulated 10000 sets of 64 $v_{\mathrm{l}}$ values. From each set of 64 $v_{\mathrm{l}}$ values we derived a $K$ based on Equation 6. With the altogether 10000 derived $K$ values, we estimated $K$ and its uncertainty for two PDFs of comparison. The calculated $\ln{K}$ (as well as their uncertainties) of the candidate PDFs with respect to the best-fit Cauchy distribution are provided in Table 8.
@@
Between the two finalist PDFs, the Cauchy distribution is preferred over the three-component normal distribution according to the $\Lambda$ standard (see Table 8). Therefore, we conclude that the sample of 64 observed $v_{\mathrm{l}}$ is well described by the simple PDF
$\frac{p}{1\,\mathrm{s\penalty 10000\ {km}^{-1}}}=\frac{1}{51.2\pi\left[1+\left(\frac{{v_{\mathrm{l}}}}{51.2\,\mathrm{km\penalty 10000\ s^{-1}}}\right)^{2}\right]}\,\,.$ (8)
@@ C.1 Millisecond pulsar age distribution
The age distribution of MSPs is poorly constrained. We approached the rough MSP age distribution by the following procedure. Firstly, we identified MSPs from the PSRCAT1 catalogue with the MSP criteria (i) and (ii) stated in Appendix A. After removing sources without constraint on characteristic age $\tau_{\mathrm{c}}$, we obtained 399 MSP $\tau_{\mathrm{c}}$ values. To derive the age distribution, we corrected the $\tau_{\mathrm{c}}$ sample by a sample of age correction factors (i.e., ratio between age and $\tau_{\mathrm{c}}$) estimated by Kiziltan10 for 42 MSPs having proper motion measurements (i.e., the reciprocals of $\tau_{c}/\tilde{\tau}^{\prime}$” reported in Table 1 of Kiziltan10). Specifically, we multiplied each of the $\tau_{\mathrm{c}}$ sample by each of the correction factor sample, which led to a sample of $399\times 42=16,758$ MSP ages. After discarding ages larger than the age of the Galaxy (i.e., 13.6 Gyr), we obtained a sample of 13,005 MSP ages (see Figure 10 for the histogram of this sample). We considered this age sample as an indicative approximation to the MSP age distribution, and adopted the age sample in the DPS analysis (see Section 4.2).
@@ C.2.1 The underlying probability density function of the observed Galactic heights
When searching for the candidates for the underlying PDF of the observed $z$ sample, we assumed that the PDF is both uni-modal and symmetric about zero. The same assumptions had been adopted in identifying the underlying PDF for the observed $v_{\mathrm{l}}$ (see Section 3). We found that the one-parameter Cauchy distribution and the one-parameter Laplace distribution are both plausible candidates for the underlying PDF of the observed $z$ (see Figure 2). However, further analysis following the method detailed in Appendix B.2 prefers the best-fit Laplace distribution over the Cauchy counterpart with $\ln{K}=7.8^{+0.7}_{-0.9}$, where $K$ is the Bayes factor. The best-fit Laplace distribution for the observed $z$ sample has the scale height $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ of 0.52 kpc. We note that $\zeta^{\mathrm{local}}_{\mathrm{obs}}$ is derived from nearby field MSP system, which should be distinguished from the scale height $\zeta_{1}$ of Galaxy-wide field MSP systems.
@@ C.2.2 The determination of the Galaxy-wide birth scale height $\zeta_{0}$
As mentioned in Section 4.1, various $\zeta_{0}$ would lead to different levels of the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$. Due to the complex distributions of the observed $z$ and $v_{\mathrm{b}}$ (see Figure 2 or Figure 14), we find it difficult to determine the $\zeta_{0}$ using statistical methods (e.g., Kolmogorov-Smirnov tests or the Wasserstein metric) that quantify the similarity between the observed samples and the DPS counterparts. On the other hand, the evolved scale height $\zeta_{\mathrm{e}}^{\mathrm{local}}$ is expected to increase with $\zeta_{0}$ (see Figure 14), and match the observed counterpart $\zeta^{\mathrm{local}}_{\mathrm{obs}}=0.52$ kpc (see Appendix C.2.1). Therefore, we reran the DPS with incrementally larger $\zeta_{0}$ until the best-fit $\zeta_{\mathrm{e}}^{\mathrm{local}}$ reaches 0.52 kpc. In this way, we found $\zeta_{0}\approx 0.32$ kpc. The dependence of the evolved $v_{\mathrm{b}}$, $z_{\mathrm{e}}$ and $v_{\mathrm{l}}$ on $\zeta_{0}$ is illustrated in Figure 14. As is shown in Figure 14, both the evolved $v_{\mathrm{b}}$ and $z_{\mathrm{e}}$ match the observed counterparts at $\zeta_{0}=0.32$ kpc. Additionally, the average levels of both evolved $|\mathrm{b}|$ and $|z_{\mathrm{e}}|$ rise with $\zeta_{0}$ as expected. In comparison, the evolved $v_{\mathrm{l}}$ stays robust with respect to changes in the birth vertical distribution of field MSP systems.
@@
Figure 14: The smoothed histogram profiles of $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ derived from dynamical population synthesis based on different Galaxy-wide birth scale height $\zeta_{0}$ of field MSP systems. Overlaid are the visualized histograms (and their smoothed profiles) of the observed $z$, $v_{\mathrm{l}}$, and $v_{\mathrm{b}}$ (see Section 2).
@@
Finally, we note that the observed $z$ sample might be biased by Galactic-latitude-related selection effects. As a major example, pulsars located at low Galactic latitudes $b$ are more susceptible to propagation effects caused by the ionized interstellar medium (IISM), which might reduce the detection rate of these low-$|b|$ pulsars. This relative drop in the number of low-$|b|$ pulsars would become increasingly significant as further-away pulsars are searched and observed with large telescopes. In Section 2.1, we found that the observed scale height of binary field MSPs is larger than solitary field MSPs, which may indicate the larger difficulty to detect binary MSPs (compared to solitary MSPs) in scattered low-$|b|$ sky regions. On the other hand, most pulsar search programs have been focusing on low-$|b|$ sky regions, which serves as another selection effect that would increase the relative number of low-$|b|$ pulsars. We note that we did not take into account the impact of complex $b$-related selection effects on the observed $z$ (and $v_{\mathrm{b}}$) in this work.
@@
Figure 16: The histogram of Galaxy-wide evolved Galactic heights $z_{\mathrm{e}}$ derived from dynamical population synthesis. Overlaid are the best-fit Laplace and Cauchy distributions of the $z_{\mathrm{e}}$ sample, with the scale parameters of 0.68 kpc and 0.48 kpc, respectively.